OpenStudy (anonymous):

What is special about the standard equation of a circle? Help please I have no idea!

OpenStudy (ankit042):

you have very weird subjective question hmm I think you can find the center and radius directly from the standard

OpenStudy (ankit042):

I don't know I am not sure but this my guess...and something I find special with circle eq lol

OpenStudy (anonymous):

I don't either. My teacher asked it and I can't move on without knowing it but I'm very confused. Equation of a circle where (h, k) is the center and r is the radius: (x − h)2 + (y − k)2 =r2 The formula for a circle (x − h)2 + (y − k)2 = r2 can also be re-arranged and written as: x minus h quantity squared all over r squared + y minus k quantity squared all over r squared = 1 so i guess you're right about finding the center and the radius. But how would I explain how you do that?

OpenStudy (ankit042):

sorry I am not sure what your question is but lets take an example (x-5)^2+(y+3)^2 = 16 can you find center and radius of this circle?

OpenStudy (anonymous):

No I have no idea :( Do I just solve that equation?

OpenStudy (ankit042):

so if you look closely you can see value of (h,k) and r hint (x-5)^2+(y-(-3))^2=(4)^2

OpenStudy (anonymous):

I have no idea I'm sorry

OpenStudy (ankit042):

k so say I ask you to write a equation of circle with radius 4 and center (5,-3) can you do that?

OpenStudy (anonymous):

So...would it be....(x-5)^2+(y+3)^2=16?

OpenStudy (anonymous):

WAIT that's what you just posted!

OpenStudy (ankit042):

Great! so can you tell me the center and radius for circle by just looking at the equation? you just have to figure out value of (h,k) and r

OpenStudy (anonymous):

so the center would be (-5,3) and r would be 16?

OpenStudy (ankit042):

Noooooooooooooooo

OpenStudy (anonymous):

Does it change after you make it an equation? Cause you said before it was (5,-3) and 4

OpenStudy (anonymous):

But in the equation it looks like (-5,3) and 16

OpenStudy (ankit042):

just check the radius again and also tell me for this new one (x-2)^2+(y+8)=36

OpenStudy (anonymous):

would it be...(2,-8) and 6?

OpenStudy (ankit042):

Yes I think you got the concept! good job :)

OpenStudy (anonymous):

Thank you!

OpenStudy (ankit042):

np :)

OpenStudy (ankit042):

And please do let me know what you teacher thinks is special in circle equation

OpenStudy (anonymous):

I guess I could say for my actual question lol is that you can find the center and the radius and that the equation is derived from Pythagorean Theorem. Not sure if it'll be right but I'll let you know lol

OpenStudy (ankit042):

idk if it helps you can add that point x and y are at a constant distance r from the center.

OpenStudy (anonymous):

Oh I'll add that too just to see if it's right, I'll let you know what she says in a little :) thanks again!

OpenStudy (ankit042):

np :) Yeah I would like to know the answer

OpenStudy (anonymous):

Okay, so basically all she wanted was that (h,k) was the center and r was the radius. Then she wanted me to solve \[(x-5)^2/16+(y+5)^2/16=1\]

OpenStudy (anonymous):

To change that into the normal circle equation and I got it right. Thank you for teaching me!!!

OpenStudy (ankit042):

haha Ok so that was special thing about circle good to know!